OpenStudy (anonymous):

Which of the following represents 3x^5/7 in radical form?

OpenStudy (anonymous):

\[3x^ \frac{ 5 }{ 7 }\]

OpenStudy (abb0t):

HINT: \(a^\frac{ P }{ x } = \sqrt[x]{a^P}\)

OpenStudy (anonymous):

so it would be \[\sqrt[7]{3x}^{5}\]

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

thanks